OpenStudy (anonymous):

what is the product of 2 sqrt 15k * 6k sqrt 3

OpenStudy (anonymous):

could you draw it out/write equation

OpenStudy (anonymous):

@kcheung1

OpenStudy (anonymous):

36k sqrt 5 k because you have 2* 6k = 12 k then sqrt 15k*3 = 12k sqrt 45 k, 45 can be broken down into 9 and 5, sqrt 9 is 3 and so you have 12k*3 = 36k and what's left inside which is sqrt 5 k

OpenStudy (anonymous):

thanks love!